The sum of square of two numbers is 34 and their product is 15. Please solve with full steps in detail.
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

The sum of square of two numbers is 34 and their product is 15.
Please solve with full steps in detail.
-----
x^2 + y^2 = 34
xy = 15
------------------------
x = 15/y
------
Substitute for "x" and solve for "y"::
(15/y)^2 + y^2 = 34
------
15^2 + y^4 = 34y^2
-----
y^4 - 34y^2 - 15^2 = 0
-------------------------------
Use the quadratic formula to solve for "y^2".
Then solve for "y".
Check for extraneous answers.
Cheers,
Stan H.
--------

Answer by Alan3354(69443)

The sum of square of two numbers is 34 and their product is 15.
Please solve with full steps in detail.
-----
x^2 + y^2 = 34
xy = 15
------------------------
x = 15/y
------
Substitute for "x" and solve for "y"::
(15/y)^2 + y^2 = 34
------
15^2 + y^4 = 34y^2
-----
y^4 - 34y^2 - 15^2 = 0
======================================== Copied from the other tutor Stanbon.
y^4 - 34y^2 + 15^2 = 0 *** not - 15^2 ***
y^4 - 34y^2 + 225 = 0
(y^2 - 25)*(y^2 - 9) = 0
--> y = 3 & 5
y = -3 & -5